OpenStudy (anonymous):

Tue’s math teacher gave 20 four-question pop-quizzes throughout the first semester. Tue remembered that she scored either 75% or 100% on all of the quizzes. If her average quiz grade for the semester is 90%, how many of each score did she earn?

OpenStudy (anonymous):

let x = the number of scores that were 75% let y = the number of scores that were 100% x + y = 20 and (0.75)x + (1.00)y = (20)(0.90) Now, you have 2 simultaneous equations in 2 variables, so just solve for x and y.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

If you have any difficulties at this point, just let me know.

OpenStudy (anonymous):

x=16 y=4

OpenStudy (anonymous):

is that close?

OpenStudy (anonymous):

Well, "close" is hard to say. It's not correct if that's what you are asking. Are you able to use substitution or elimination?

OpenStudy (anonymous):

This particular problem lends itself very easily to elimination. If you rearrange the first equation, you get: y = 20 - x If you put that into equation #2 you get: (0.75)x + (1.00)(20 - x) = (20)(0.90) which gives easy numbers to work with. Now you have one equation in one variable "x". Can you solve for "x"?

OpenStudy (anonymous):

i tried that and got the work right. But x came out as 16....again

OpenStudy (anonymous):

x=0.8?

OpenStudy (anonymous):

np, I'll show you what I did: (0.75)x + (1.00)(20 - x) = (20)(0.90) -> (0.75)x + 20 - x = 18 (0.25)x = 2 -> x = 8 So, y = 12

OpenStudy (anonymous):

wouldn't it be -.25?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

ok thanks for your help

OpenStudy (anonymous):

You're welcome!